### v3

# Fine-grained Stateful Knowledge Exploration: Effective and Efficient Graph Retrieval with Large Language Models

Abstract
Large Language Models (LLMs) have shown impressive capabilities, yet updating their knowledge remains a significant challenge, often leading to outdated or inaccurate responses. A proposed solution is the integration of external knowledge bases, such as knowledge graphs, with LLMs. Most existing methods use a paradigm that treats the question as the objective, with relevant knowledge being incrementally retrieved from the knowledge graph. However, this strategy frequently experiences an mismatch in the granularity of knowledge between the target question and the entities and relations being retrieved. As a result, the information in the question cannot precisely correspond to the retrieved knowledge. This may cause redundant exploration or omission of vital knowledge, thereby leading to enhanced computational consumption and reduced retrieval accuracy. In this paper, we propose a novel paradigm of fine-grained stateful knowledge exploration, which addresses the ‘information granularity mismatch’ issue. We extract fine-grained information from questions and explore the semantic mapping between this information and the knowledge in graph. By dynamically updating the mapping records, we avoid redundant exploration and ensure no pertinent information is overlooked, thereby reducing computational overhead and improving the accuracy of knowledge exploration. The use of fine-grained information also eliminates the need for a priori knowledge, a common requirement in existing methods. Experiments on multiple datasets revealed that our paradigm surpasses current advanced methods in knowledge retrieval while significantly reducing the average number of LLM invocations. The code for this paper can be found at https://github.com/nnnoidea/stateful-KGQA.
1 Introduction
Large language models (LLMs) have demonstrated considerable efficacy in a range of natural language processing tasks, as evidenced by recent studies Achiam et al. (2023); Cheng et al. (2023); Meta (2024); Zhao et al. (2024). Nevertheless, there are certain limitations to be considered when employing LLMs in question-answering tasks. The knowledge of LLMs remains static, fixed at the point of their training, which presents a challenge in incorporating timely updates. Moreover, in response to queries with which they are unfamiliar, LLMs may occasionally provide incorrect answers accompanied by an overconfident tone Luo et al. (2023), a phenomenon that is often referred to as hallucination Bang et al. (2023); Ji et al. (2023). In light of these challenges, an effective solution is the integration of external knowledge, specifically knowledge graphs (KGs), to enhance the question-answering proficiency of LLMs.
Figure 1: A simple example illustrating the differences between the paradigm we propose and the existing main stream paradigm. Stateless exploration may perform redundant exploration at times and may miss information. In contrast, the stateful exploration we propose can avoid redundancy and omissions.
A number of studies have been conducted with the objective of combining external knowledge with LLMs. One straightforward approach is to prompt LLMs to directly generate the corresponding knowledge base operation statements based on questions. This method has been demonstrated to be effective in various studies, as referenced in Hu et al. (2023); Wang et al. (2023). Additionally, chain-of-thought Wei et al. (2022b) is employed to assist LLMs in decomposing input questions into multiple sub-questions Li et al. (2023a); Zhao et al. (2023). However, these methods do not take into account the awareness of the nature of external knowledge sources on the part of LLMs. Consequently, the instructions or sub-questions generated by LLMs may not be aligned with the external knowledge. To avoid the mismatch issue due to unawareness of the knowledge base, current methods no longer try to get the entire knowledge path in a single step. Instead, they acquire it incrementally through iteration. These methods furnish the LLM with subgraph information from the knowledge base, thereby averting the occurrence of mismatch during each iteration, such as Sun et al. (2024).
However, the existing methods only avoid the mismatch at the character level without noticing the mismatch at the information granularity. The question, as a sentence containing multiple pieces of information, has an information granularity that is significantly larger than the knowledge granularity of entities and relations in the knowledge graph. This leads to two problems. First, the information in question does not correspond precisely to the retrieved knowledge, so existing methods have to engage in stateless iterative exploration. In stateless iterative exploration, the same sentence is matched with subgraphs containing different entities and relations in different rounds. It is important to note that some information in the sentence may be matched repeatedly in different rounds, while other information may be ignored from the beginning to the end. An example is shown in the upper half of the Fig. 1. On the one hand, repeated matching increases the number of exploration rounds, resulting in higher computational consumption, especially for large language models. On the other hand, if some information is ignored, the retrieved knowledge will be incomplete and the question-answering effect will be compromised. Second, it is evident that current methodologies require the artificial provision of fine-grained topic-entity as a starting point for knowledge exploration paths Sun et al. (2024); Chen et al. (2024). This will undoubtedly result in additional manual effort in practice. Some approaches are vaguely aware of the information granularity mismatch problem and split the query into multiple clauses Chen et al. (2024), but do not fundamentally address the mismatch in information granularity.
To solve the above problems caused by the mismatch of information granularity, in this paper we propose a new paradigm of fine-graind stateful knowledge exploration, as shown in the lower part of Fig. 1. The new paradigm extracts fine-grained, sufficient information from the question to capture most of its semantics. It also transforms the goal of knowledge exploration from coarse-grained to multiple independent fine-grained information. We call this fine-grained information clues. When performing knowledge exploration, the knowledge corresponding to a clue is explored iteratively from the knowledge graph. We call this correspondence between the clues and the knowledge in the graph a mapping. In this stateful knowledge exploration process, keeping track of the clues and their mappings ensures that different rounds of exploration do not explore duplicate information, and also avoids ignoring some information. An example of this is shown in Figure 1. In this new paradigm, using clues as the starting point of knowledge exploration eliminates the need for human-provided topic entities, a common requirement in existing methods.
The main contributions of this paper are listed as follows:
  • We point out the issue of information granularity mismatch in the existing mainstream paradigms, which forces current methods to perform stateless iterative exploration. This results in increased computational consumption in the knowledge exploration process, incomplete retrieval of knowledge, and additional requirements for prior knowledge.
  • We propose a new paradigm for stateful knowledge exploration based on fine-grained information, which resolves the issue of information granularity mismatch. The new paradigm eliminates the need for prior knowledge. Stateful iterative exploration significantly reduces the computational consumption in the knowledge retrieval process and improves the effectiveness of knowledge retrieval.
  • We conducted experiments on multiple datasets. The results show that our proposed new paradigm performs excellently on all datasets. The new paradigm not only outperforms existing methods in knowledge graph question-answering but also significantly reduces the number of LLM calls, sometimes to only a fraction of that required by existing methods.
2 Related work
2.1 Document Retrieval Enhanced Models
One direct approach to addressing the generation of illusions by LLMs is the use of external knowledge Mialon et al. (2023) to enhance the model’s ability to perceive the real world. This has led to the emergence of numerous knowledge-based models. Information retrieval is commonly employed in developing knowledge-based chatbots Shuster et al. (2021). BlenderBot2 Xu et al. (2021) integrates Internet search capabilities. SeeKeR Shuster et al. (2022a) surpasses BlenderBot2 Xu et al. (2021) by executing three modular tasks using a single language model: generating search queries, extracting relevant knowledge from retrieved documents, and generating the final response. BlenderBot3 Shuster et al. (2022b) fine-tunes a 175-billion-parameter OPT Zhang et al. (2022) on a combination of 20 question-answering and dialogue datasets. Atlas Izacard et al. (2023) stands out as a top model on the KILT benchmark Petroni et al. (2021), which includes 11 knowledge-oriented tasks, including Wizard of Wikipedia Dinan et al. (2019).
2.2 Knowledge base Enhanced Models
There are diverse ways to enhance models based on knowledge, aiming to address issues arising from complex reasoning, hallucination, and poor interpretability. Cao et al Cao et al. (2021) improved implicit event causality extraction by leveraging structured knowledge bases. Zhao et al Zhao et al. (2023) recently proposed a Verification and Editing (VE) framework. This framework improves the Chain of Thought (CoT) reasoning in large language models (LLMs) by integrating retrieval systems Wei et al. (2022a). Researchers have also introduced the Chain-of-Knowledge (CoK) Li et al. (2023b), a novel framework designed to enhance the factual correctness of LLMs using heterogeneous knowledge sources. This ensures that inaccuracies in the foundational principles do not propagate to subsequent steps. Think-on-Graph(ToG) Sun et al. (2024) is a plug-and-play framework that seamlessly applies to various LLMs and KGs. ToG also enhances the reasoning capabilities, enabling smaller LLMs to compete with larger LLMs.
3 Methodology
3.1 Problem definition
Knowledge Graph Question-Answering is the task of retrieving relevant knowledge from a knowledge graph to answer given questions. In our proposed new paradigm, we first obtain a set of clues $clue_{all}=\{\text{clue}_{1},\text{clue}_{2},\dots\}$ from the question q. For each clue in the $clue_{all}$, we explore its corresponding relations $r$ and entities $e$ in the knowledge graph, forming a knowledge path $P$ = $\{(e^{0},r^{1},e^{1}),(e^{1},r^{2},e^{2}),\dots,(e^{i-1},r^{i},e^{i})\}$. This path then serves as a base for the large language model to answer the question.
3.2 Fine-grained clues extraction
To achieve stateful knowledge exploration, we need to obtain fine-grained information, which we refer to as clues. Clues should not only characterize the information of the question, but also match the granularity of information, or information content, with the entities and relations in the knowledge graph. Intuitively, we posit that the majority of information (not all information) within a question can be characterized by keywords originating from the question. Simultaneously, the granularity of keywords is comparable to that of entities in the knowledge graph. The use of these keywords as clues for knowledge retrieval helps in addressing the given question. For example, in the question ‘What is the population of the capital of China?’, the key information required is ‘population’, while ‘China’ and ‘capital’ serve to narrow the context. (China, capital, population) are the keywords in the given question. However, it is important to note that different knowledge graphs have different ontology designs. The same clue may correspond to an entity in graph A while being related to a relation in graph B. To heighten the universality of our approach, we refrain from making explicit distinctions between clues that map to entities or relations during the clue extraction phase. Instead, we simply prompt the LLM to extract the entire set of clues.
3.3 Starting point
After obtaining the fine-grained clues, we first determine the starting point of the exploration. Unlike existing methods that require human-provided topic entities as starting points, the clues we extract can serve directly as starting points. It was mentioned above that some clues may map to relations, but information extracted from a semantically complete sentence necessarily has one or more mappings to entities, such as ‘China’ in the example above. Sentences without entities cannot effectively express semantics. This is not only consistent with our knowledge, but also verified in our experiments. We retrieve entities that match the clue characters in the knowledge graph and use any successful matching entity as a starting point. Then, we will gradually explore the entire knowledge path based on this starting point.
3.4 Stateful knowledge path exploration
Through iterative exploration, we explore the entities or relations mapping to each clue, forming a knowledge path. Each round of mapping exploration consists of two steps: relation-based mapping exploration and entity-based mapping exploration. For better understanding, we describe the process using the $n$-th iteration as an example. In the previous $n-1$ iterations, we have obtained $n-1$ triples $(e^{i-1},r^{i},e^{i})$, where $e^{i-1}$ is the tail entity of the $(i-1)$-th triple and the head entity of the $i$-th triple, and $r^{i}$ is the relation of the $i$-th triple. The starting point $e^{0}$ is the head entity of the first triple. The set of clues already explored is $C$, and the remaining set of clues $C^{\prime}=clue_{all}-C$.
Relation-Based Mapping Exploration
The first step in each round of mapping exploration is relational mapping exploration. In the $n$-th iteration, we first obtain the relations $R=\{r_{1},r_{2},r_{3},r_{4},...\}$ associated with the entity $e_{n-1}$ from the graph, where $e^{n-1}$ is the tail entity of the triple obtained in the $(n-1)$-th exploration. We provide $R$ and $C^{\prime}$ to LLM and ask LLM to determine if there exists a $clue_{i}\in C^{\prime}$ that maps to any $r$. The mapping result is denoted as:
[…]
If r-mapping is empty, the exploration terminates.
Entity-Based Mapping Exploration
[…]
Three different results can occur:
[…]
3. e-mapping is empty and then terminate the exploration.
3.5 Ending exploration and answering questions
Every step in the exploration process might result in multiple outcomes, thus generating multiple branches. To prevent an explosion in the number of branches, existing methods opt to select the top-N branches as final results. However, this approach could inevitably lead to omissions, particularly when the number of correct results exceeds N. In our method, we do not filter branches based on the quantity, but instead, on whether all clues can be mapped. Therefore, during the exploration process, any branch unable to find the next clue mapping will be terminated. If a branch that finds all clue mappings emerges, then the exploration of all other branches concludes simultaneously. At the end of the exploration, if there exists a knowledge path where all clue mappings have been found, it is offered to the large language model to answer the question. If no such path exists, then it is assumed that the exploration failed, and the question is directly posed to the large language model for an answer.
Table 1: Results for MOOC Q&A.
Method 1-Hop True/False Multi-Hop True/False 1-Hop Query
partial match complete match
without external knowledge
IO prompt Brown et al. (2020) w/ChatGPT 8.0 7.3 10.0 7.5
CoT Wei et al. (2022b) w/ChatGPT 3.4 2.2 10.0 6.7
with external knowledge
ToG Sun et al. (2024) w/GLM4-9B-Chat 48.7 37.6 24.5 8.2
ToG Sun et al. (2024) w/ChatGPT 53.8 39.2 65.6 46.9
ToG Sun et al. (2024) w/GPT-4 54.3 41.7 82.6 56.5
Ours w/GLM4-9B-Chat 84.1 55.0 77.1 67.9
Ours w/ChatGPT 89.1 71.1 83.1 76.4
Ours w/GPT-4 94.4 74.4 84.2 77.8
4 Experiments
4.1 Experimental Settings
Datasets and Evaluation Metrics. To evaluate the performance of our proposed paradim, We selected two open-source knowledge graphs and one self-constructed graph as external knowledge bases for experimentation. The open-source knowledge graphs are MOOC Q&A Yu et al. (2020) and Freebase Bollacker et al. (2008). The graph that we constructed ourselves is not yet publicly available and is referred to in this paper as the agricultural knowledge graph.
MOOC Q&A is a dataset collected from MOOC websites, comprising a knowledge graph and associated question-answer sets. The knowledge graph includes 7 categories of entities, totaling 15,388, with 11 types of relationships forming 277,487 triplets. The question-answer sets contain True/False Question and Query Question.
Freebase is a large-scale, semi-structured database supported by Google, designed to collect and connect information about millions of entities and their relationships worldwide. It covers a wide range of topics such as people, places, books, movies, music, etc., with each entity having a unique identifier. We conduct experiments on three QA sets with Freebase as external knowledge base: WebQSP Yih et al. (2016) and WebQuestions Berant et al. (2013).
The agricultural knowledge graph that we constructed includes over 100,000 entities and 1 million triples, on which basis we built single-hop and multi-hop question-answering sets.
[…]
Baselines. We compared our approach with three baseline methods: standard prompting (IO prompt) Brown et al. (2020), chain of thought prompting (CoT prompt) Wei et al. (2022b) and ToG Sun et al. (2024). The knowledge base and Q&A sets in MOOC Q&A are highly interconnected, with most questions requiring knowledge retrieval from the knowledge base for answers. IO prompt and CoT prompt are two methods without external knowledge, and they are used to illustrate the proportion of questions requiring knowledge. ToG is a previous state-of-the-art method in knowledge base question-answering, used to validate the effectiveness of our knowledge retrieval approach.
Experiment Details. To ensure the reliability and reproducibility of the experiments, we set the temperature parameter to 0 for all LLMs. The two hyperparameters of ToG are set to their default values. We modified all prompts in ToG to be suitable for MOOC Q&A and the agricultural knowledge grpah. During the experiment, we observed that ToG would invoke LLM hundreds of times to answer certain questions, exceeding ten minutes. To mitigate the impact of these extreme cases, we terminated the exploration for all methods when it exceeded 30 LLM calls.
Table 2: Results for WebQuestions and WebQSP. To better illustrate whether these methods answer questions based on the inherent knowledge of LLMs or the retrieved knowledge, we merged the responses of these methods with the IO answers of LLMs and conducted additional comparisons. These are respectively represented as ToG + IO and Ours + IO. For each dataset, we conducted experiments on 300 data samples.
Method WebQuestions WebQSP
partial match complete match partial match complete match
without external knowledge
IO prompt Brown et al. (2020) w/Llama3-8B-Instruct 44.4 37.2 52.4 27.0
IO prompt Brown et al. (2020) w/GLM4-9B-Chat 52.4 33.4 60.1 31.8
COT Wei et al. (2022b) w/Llama3-8B-Instruct 54.3 35.8 62.1 33.7
COT Wei et al. (2022b) w/GLM4-9B-Chat 54.8 34.3 63.0 32.5
with external knowledge
ToG Sun et al. (2024) w/Llama3-8B-Instruct 55.1 36.7 55.6 32.3
ToG Sun et al. (2024) w/GLM4-9B-Chat 63.3 37.8 63.3 30.6
ToG + IO w/Llama3-8B-Instruct 67.3 43.9 69.7 39.4
ToG + IO w/GLM4-9B-Chat 71.4 46.9 74.0 38.8
Ours w/Llama3-8B-Instruct 70.3 44.6 70.8 40.4
Ours w/GLM4-9B-Chat 66.3 38.8 62.5 36.4
Ours + IO w/Llama3-8B-Instruct 81.2 50.5 75.0 43.3
Ours + IO w/GLM4-9B-Chat 77.9 50.2 73.9 43.2
Table 3: Results for the agricultural graph. The knowledge in this graph and the corresponding question-answering sets are relatively specialized.
Method 1-Hop Query multi-Hop Query
partial match complete match partial match complete match
ToG Sun et al. (2024) w/GLM4-9B-Chat 42.9 30.3 10.0 8.2
Ours w/GLM4-9B-Chat 48.2 32.1 9.8 9.3
4.2 Performance Comparison
We conducted experiments on three question-answering sets of MOOC Q&A: 1-hop True/False, multi-hop True/False, and 1-hop Query. The results are shown in Table 1. It is evident that the baseline methods, IO and CoT, which do not utilize external knowledge, exhibit very low accuracy rates in answering questions. This implies that these questions are not familiar to LLMs, and answering them requires a high degree of reliance on knowledge within the knowledge graph.
On the True/False sets, our method demonstrates outstanding performance, surpassing the previous state-of-the-art method ToG. Even the performance of our method based on GLM4-9B surpasses that of ToG based on GPT-4. This demonstrates the significant advantages of the new paradigm we propose over existing paradigms. The method we proposed showed improvements when different models such as GLM4-9B, ChatGPT, and GPT4 were used. This indicates that our method is very friendly to LLMs with fewer parameters, enabling better results at a lower cost, while still offering considerable room for improvement with better models.
On the Query set, considering the partial match metric, in the case of GPT4, ToG’s results are only slightly worse than our method. However, when based on ChatGPT and GLM4-9B, ToG performs significantly worse than our method. This once again highlights the usability of our approach with LLMs that have smaller parameter sizes. Examining the complete match metric, our method consistently outperforms ToG across all scenarios. This indicates that while ToG can identify partial answers to questions, it struggles to comprehensively retrieve all possible answers to a given question. In contrast, our method excels in providing a comprehensive set of answers to the posed questions.
The results of WebQuestions and WebQSP are shown in Table 2. Due to economic considerations and OpenAI’s policies, we did not test ChatGPT and GPT-4 for WebQSP and WebQuestions. From the baseline, we can observe that LLMs can answer a considerable portion of the questions solely based on their own knowledge, indicating that these questions are relatively familiar to LLMs. To better illustrate whether these methods answer questions based on the inherent knowledge of LLMs or the retrieved knowledge, we merged the responses of these methods with the IO answers of LLMs and conducted additional comparisons. These are respectively represented as ToG + IO and Ours + IO. It can be seen that even when using a massive external knowledge base like Freebase, our method still maintains a comprehensive advantage over ToG. It is worth noting that ToG requires manual provision of a topic entity, which means it can directly find a starting point in the knowledge base. Our method does not use manually provided topic entities. Comparing the result improvements of both methods after adding IO, it can be seen that IO provides a greater boost to the results of our method. This means that the results of our method are more often based on correctly retrieving knowledge from the knowledge base and answering questions, rather than relying on the inherent knowledge of the LLM.
The results of the agricultural graph are shown in Table 3. It can be seen that the method we proposed outperforms ToG on single-hop questions, but does not show an advantage on multi-hop questions. This agricultural dataset, which we constructed ourselves, has highly specialized questions, and multi-hop question-answering is relatively more complex. This may indicate that our method is not superior to existing methods when dealing with specialized and complex questions.
In summary, the new method we proposed has obvious advantages over existing methods, with significantly higher accuracy and completeness of answers compared to existing methods. However, when encountering overly specialized and complex questions, our method performs on par with existing methods.
4.3 Computational Cost
In the knowledge base question-answering process, the majority of computational resources are dedicated to utilizing the Large Language Model, while the remaining computational consumption is relatively negligible. We calculate the average number of LLM calls during the question-answering process to illustrate the computational resource consumption. The results are shown in Table 4 and 5.
The results show that our method and ToG exhibit consistency in the number of LLM invocations across different question-answering sets. On more complex question sets, there are more LLM invocations. When encountering the same question-answering set, our method invokes the LLM significantly fewer times than ToG, roughly only one-third to half as many. We believe that this advantage is due to the stateful exploration we proposed, which can avoid redundant exploration.
Table 4: Average Number of Calls to LLMs On MOOC Q&A
Method 1-Hop True/False Multi-Hop True/False 1-Hop Query
ToG Sun et al. (2024) w/ChatGLM4-9B 14.76 16.82 8.94
ToG Sun et al. (2024) w/ChatGPT 19.81 22.16 11.2
ToG Sun et al. (2024) w/GPT-4 22.6 26.17 8.06
Ours w/ChatGLM4-9B 6.64 9.32 4.71
Ours w/ChatGPT 5.86 8.18 5.58
Ours w/GPT-4 5.19 6.93 4.97
Table 5: Average Number of Calls to LLMs On the agricultural graph
Method 1-Hop Query Multi-Hop Query
ToG Sun et al. (2024) w/ChatGLM4-9B 11.64 16.2
Ours w/ChatGLM4-9B 5.72 9.45
Table 6: Ablation Studies on WebQuestions set. We conducted three experiments: removing the stateful exploration, removing the ending flag, and removing both.
Method partial match complete match avg LLM Calls
Ours 70.3 44.6 8.76
w/o Stateful exploration 65.7 44.3 9.06
w/o ending flag 59.4 42.6 9.12
w/o both 51.4 38.7 15.9
4.4 Ablation Studies
We conducted ablation studies and the results of ablation studies are shown in Table 6. We conducted three experiments: removing the stateful exploration, removing the termination flag, and removing both. Removing the ending flag also means that it is impossible to determine whether knowledge should be provided to the LLM or whether to directly ask the LLM a question. It can be seen that both stateful exploration and the ending flag contribute to reducing the number of LLM invocations, and when both parts are removed, the number of LLM invocations increases significantly. This is because both steps are based on fine-grained knowledge. Fine-grained knowledge, by mapping to the retrieved knowledge, naturally avoids redundant exploration.
When either of the two steps was removed, the question-answering results declined, with a more significant drop observed when the ending flag was removed. We analyzed that this is because the ending flag also serves as the basis for determining whether to provide knowledge to the LLM. After removing the ending flag, it becomes impossible to judge whether the retrieval is correct and complete, so the retrieved knowledge must be provided to the LLM every time. When retrieval fails, the erroneously retrieved knowledge interferes with the LLM’s response, leading to a decline in results.
5 Conclusion
In this paper, we propose a new paradigm of fine-grained stateful exploration. It aims to address the widespread issue of knowledge granularity mismatch in existing knowledge graph question-answering paradigms. We analyze that this problem leads to redundant exploration or omission of important information during the knowledge exploration process, ultimately resulting in increased computational consumption and degraded question-answering performance. In the new paradigm we propose, fine-grained information is mapped to the retrieved knowledge, effectively avoiding repetition and omissions. We conducted experiments based on three different knowledge graphs and corresponding multiple question-answering sets, and the results validated our design concepts. Our method significantly outperforms existing methods in terms of both question-answering results and computational efficiency. Additionally, the experiments demonstrate that our method still performs well with LLMs that have fewer parameters, showcasing great practical potential. Ablation studies also prove the effectiveness of each component of our design.